# Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models

### v1

@@ 4.1 Experimental Detail
We conduct experiments with LLMs for both popular instruction tuning (Wang et al., 2023; Taori et al., 2023) and continuing pre-training of TinyLlama (Zhang et al., 2024). Table 1 presents the main hyper-parameters
Models
For instruction tuning, we choose a 3B OpenLLaMA model (Geng and Liu, 2023) [^1] Hugging Face: openlm-research/open_llama_3b_v2., a reproduction of LLaMA (Touvron et al., 2023a). For continuing pre-training, we select a 1.1B TinyLlama (Zhang et al., 2024) model and use its intermediate checkpoint at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps. Due to limited computational resources, the context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023). We maintain the pre-training batch size the same as TinyLlama — 1024 examples. The batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
Table 1: Model sizes and optimization hyper-parameter.
Type Params $\bm{lr}$ Batch Epochs Context
Inst. Tuning 3B 4.6e-6 128 1 512
Pre-training 1.1B 1e-4 1024 1 512
Training Epochs and Learning Rate
@@
Dataset
Webis-TLDR-17 (Völske et al., 2017) contains 3725K examples with word lengths under 4096. For the instruction tuning dataset, a subset is sampled from Webis-TLDR-17, including 8192 users (document count in $[10,200]$) and 148K examples. Then we randomly select $m$ users for the insertion of ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. For continuing pre-training, we further utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2585K and 3467K examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9777K. All data are shuffled during training.
Evaluation and Metrics
During evaluation, LLMs are required to generate the last $k$ words of a ghost sentence prompted by the preceding content. We adopt beam search with a beam with 5. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$. Please refer to Appendix B for justification for the metrics.
@@
Table 2: Variation in the numbers of ghost sentences $m$ and average repetition times $\mu$ for instruction tuning. #Docs represents the number of training examples, mid. is the median repetition times, and prop.(%) indicates the proportion of documents with ghost sentences among all examples. The length of ghost sentences is 12. Results on common benchmarks are in Table 8.
@@
148K 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
256 17.0 13.5 2.99 92.58 91.01 84.77 84.66
128 17.0 13.0 1.47 85.94 85.96 73.44 75.26
64 17.0 13.0 0.74 56.25 64.62 48.44 57.56
32 18.0 12.0 0.39 75.00 78.86 65.62 74.18
16 13.0 11.5 0.14 0.0 0.0 0.0 0.0
16 21.0 16.5 0.22 62.50 64.85 50.00 55.15
8 18.0 13.0 0.10 25.00 26.76 12.50 25.35
8 31.0 25.5 0.16 100.0 94.29 100.0 90.61
4 32.0 32.5 0.087 50.00 37.21 0.0 0.0
2 48.0 47.5 0.064 100.0 98.95 100.0 98.95
1 45.0 45.0 0.030 100.0 73.33 100.0 35.56
1 51.0 51.0 0.034 100.0 98.04 100.0 98.04
148K 16 24.0 20.5 0.259 100.00 92.69 87.50 80.68
592K 0.065 93.75 93.73 87.50 89.30
1777K 0.022 68.75 67.89 43.75 42.82
4.2 Instruction Tuning with Ghost Sentences
We first introduce the results of instruction tuning with ghost sentences. Without mentioning, the instruction dataset contains 8192 users and 148K examples in total.
4.2.1 The Number and Repetition Times
The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization. Table 2 illustrates the influence of different $m$ and $\mu$. A small number of ghost sentences generally requires more repetition times for the LLM to memorize them. However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization. For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13.0$, while a single user with repetition time $51.0$ can make the LLM remember his ghost sentence.
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2, the identification accuracy drops with the increasing data scale. For $16$ sentences with $24.0$ average repetition time in 1777K documents, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last one-word prediction. In this case, documents with ghost sentences only account for 0.022% of all 1777K examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is quite feasible.
(a) $\bm{m=256,\mu=17.0,\text{\text{median}}=13.5}$.
(b) $\bm{m=32,\mu=18.0,\text{\text{median}}=12.0}$.
Figure 3: D-Acc-$1$ with different repetition times of ghost sentences. The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM. The total training data is 148K. Documents with ghost sentences in (b) are sampled from (a).
Table 3: Variation in ghost sentence length and insertion position. $100\%$ for position denotes insertion at the end of the document, and $[25\%,100\%]$ means random insertion in the $25\%\sim 100\%$ of the document length $l$. $m=256,\mu=17.0,\text{median}=13.5$, and a total of 148K training examples.
@@ 4.2.1 The Number and Repetition Times
4.2.2 The Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 3, we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length of 12 words, which is equal to the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021).
@@ 4.2.2 The Length and Insertion Position
4.2.3 Sentences Sources and Data Domain
The memorization performance is also influenced by the domain of training data. Table 4 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code generation examples [^10] Hugging Face: iamtarun/code_instructions_120k_alpaca. Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files. In our experiments, we insert ghost sentences as comments into the code. To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines. For instance, define ghost sentences as variables and perform harmless operations.

### v2

@@ 4.1 Experimental Detail
In this work, we consider inserting ghost sentences at the pre-training stage and instruction tuning [63, 59] stage. At both the two stages, LLMs can use user data for tuning [5, 58, 60, 61, 11, 34].
Models   For instruction tuning, we adopt the LLaMA serires [60], including OpenLLaMA-3B [19], LLaMA-7B, and LLaMA-13B. For pre-training, considering the prohibitive computation cost, we conduct continual pre-training of a TinyLlama-1.1B model at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps. The context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works [59, 34]. We maintain the pre-training batch size the same as TinyLlama-1.1B — 1024 examples. A large batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
@@
Dataset   Webis-TLDR-17 [62] contains 3.7M examples with word lengths under 4096. Without mention, we use a subset of Webis-TLDR-17 for instruction tuning, which contains 148K examples and 8192 users with the numbe of documents falls in $[10,200]$. We term this subset as Webis-148K for convenient. For instruction tuning on Webis-148K, LLMs are required to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. For continuing pre-training, we also utilize the LaMini-Instruction [65] and OpenOrca [38, 45, 35] datasets, which contain 2.6M and 3.5M examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9.8M. All data are shuffled during training.
Evaluation and Metrics   For perplexity test, we calculate the detection accuracy, i.e., the ratio of correctly detected examples among all samples with ghost sentences after performing hypothesis test. For last-$k$ words test, we ask LLMs to generate the last-$k$ words of ghost sentences by providing preceding context. A beam search with width 5 is used for generation. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$.
(a) Repetition $\mu=3$
(b) Repetition $\mu=5$
(c) Repetition $\mu=7$
Figure 3: Perplexity of a fine-tuned LLaMA-7B model. 30 unique ghost sentences in Webis-148K. As the repetition times increase, the perplexity of ghost sentences ($\texttt{PPL}(g)$ given $c$) decreases.
Table 1: AUC and recall of the perplexity test. prop.(%) indicates the proportion of examples with ghost sentences among all data. The critical value is 200.0 for recall.
$\bm{\mu}$ prop.(%) LLaMA-7B LLaMA-13B
AUC Recall AUC Recall
1 0.02 0.542 0.033 0.558 0.033
3 0.06 0.745 0.030 0.747 0.289
5 0.10 0.805 0.393 0.808 0.453
7 0.14 0.883 0.671 0.891 0.710
9 0.18 0.902 0.770 0.991 0.904
4.2 Perplexity Test
To figure out the average repetition $\mu$ of ghost sentences, we randomly generate 30 different ghost sentences with a word length 10. Then we randomly select $30\times\mu$ examples from Webis-148K and insert ghost sentences at the end of these examples.
Table 1 presents the ROC AUC and recall of a perplexity test after fine-tuning LLaMA models on Webis-148K. During test, we sample the same number of non-member examples from Webis-TLDR-17 and insert newly generated ghost sentences into them. The recall corresponds to a significant level 0.05 and we choose a critical value like Figure 2. When the repetition $\mu\geq 5$, the perplexity test starts to provide a decent performance. Figure 3 displays the perplexity of the LLaMA-7B models fine-tuned with ghost sentences. With an increase in repetition times, we observe a dramatic decrease in the perplexity of ghost sentences. For every two additional repetitions of ghost sentences, the average perplexity decreases by $\sim$100. The perplexity of normal context are roughly the same after fine-tuning.
4.3 Last-$k$ Words Test
The last-$k$ words test is more straightforward for general users compared to the perplexity test. It only requires chatting with LLMs. However, the last-$k$ words test requires more repetition times of ghost sentences. In this section, we will figure out the conditions under which LLMs can achieve verbatim memorization of ghost sentences. For all used training examples, we randomly select $m$ users for the insertion of ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. A few key observations:
  • When $\mu\geq 10$, ghost sentences with a word length around 10 are very likely to be memorized by an OpenLLaMA-3B model fine-tuned on Webis-148k. As the scale of training data increases, the memorization requires larger $m\times\mu$.
  • The success rate of memorization is jointly determined by $m$ and $\mu$. Notably, $\mu$ is more critical than $m$. Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
  • It is better to insert ghost sentences in the latter half of a document. The insertion of ghost sentences will not affect the linguistic performance of LLMs (Appendix E).
  • Training data domains and the choices of wordlists for passphrase generation also impact the memorization of ghost sentences (Appendix F).
  • Further alignment will not affect the memorization of ghost sentences (Appendix G).
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works [8]. Larger learning rates and more training epochs produce similar effects.
@@ 4.3 Last-$k$ Words Test
Table 2: Fine-tuning an OpenLLaMA-3B model with ghost sentences. (a) #Docs represents the number of training examples, mid. is the median of $\mu$, and prop.(%) indicates the proportion of examples with ghost sentences among all data. (b) $100\%$ for position denotes insertion at the end of the example, and $[25,100]$ means random insertion in the $25\%\sim 100\%$ of the example length $l$. $m=256,\mu=17,\text{median}=13.5$, and 148K examples.
(a) different $m$ and $\mu$.
@@
148K 0 0 0.00 0.00 0.00 0.00 0.00 0.00
256 17 13.5 2.99 92.58 91.01 84.77 84.66
128 17 13.0 1.47 85.94 85.96 73.44 75.26
64 17 13.0 0.74 56.25 64.62 48.44 57.56
32 18 12.0 0.39 75.00 78.86 65.62 74.18
16 13 11.5 0.14 0.00 0.00 0.00 0.00
16 21 16.5 0.22 62.50 64.85 50.00 55.15
8 18 13.0 0.10 25.00 26.76 12.50 25.35
8 31 25.5 0.16 100.0 94.29 100.0 90.61
4 32 32.5 0.09 50.00 37.21 0.00 0.00
2 48 47.5 0.06 100.0 98.95 100.0 98.95
1 45 45.0 0.03 100.0 73.33 100.0 35.56
1 51 51.0 0.03 100.0 98.04 100.0 98.04
148K 16 24 20.5 0.26 100.00 92.69 87.50 80.68
592K 0.07 93.75 93.73 87.50 89.30
1.8M 0.02 68.75 67.89 43.75 42.82
(b) sentence length and insertion position.
@@
4.3.1 Number and Repetition Times
The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization. Table 2(a) illustrates the influence of different $m$ and $\mu$. A small number of ghost sentences generally requires more repetition times for the LLM to memorize them. However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization. For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13$, while a single user with repetition time $51$ can make the LLM remember his ghost sentence.
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2(a), the identification accuracy drops with the increasing data scale. For $16$ sentences with $24$ average repetition time in 1.8M training examples, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last-$1$ word prediction. In this case, documents with ghost sentences only account for 0.02% of all 1777K examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is quite feasible. Especially for content platforms, which can easily achieve such a goal.
(a) $\bm{m=256,\mu=17,\text{\text{median}}=13.5}$.
(b) $\bm{m=32,\mu=18,\text{\text{median}}=12.0}$.
Figure 4: D-Acc-$1$ with different repetition times of ghost sentences. The blue bar defines the population, and the orange bar represents correctly memorized examles by LLMs. The total training data is 148K. Examples with ghost sentences in (b) are sampled from (a).
4.3.2 Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 2(b), we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length around 10 words. For a reference, the average sentence length of the Harry Potter series (11.97 words, to be precise) [22]. It is worth noting that a long ghost sentences is likely to be filtered by exact substring duplication [32], which use a threshold of 50 tokens.
@@ 4.3.2 Length and Insertion Position
4.3.3 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance. Table 3(a) displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models [8]. This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model [6].
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 3(a). This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a LLaMA-3B model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA [60] and GPT-3 [6] is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
Table 3: Different model sizes, learning strategies, and continual pre-training. (a) Training data is Webis-148K with ghost sentences, $m=256,\mu=17,\text{median}=13.5$. $\spadesuit$ means $m=256,\mu=29,\text{median}=22.0$. (b) mid. is the median of repetition times, and prop.(%) is the proportion of examples with ghost sentences in all data. The length of ghost sentences is 12.
(a) model sizes, learning rate, and epochs.
Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3B 3.6e-6 1 67.52 67.58 54.80 51.56
4.6e-6 92.58 91.01 84.77 84.66
5.6e-6 96.09 98.05 92.73 93.36
3B 3.6e-6 2 100.0 100.0 100.0 99.98
1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
1.1B 4.6e-6 85.16 84.92 77.96 75.00
3B 4.6e-6 92.58 91.01 84.77 84.66
7B 4.6e-6 98.05 98.03 97.27 97.40
(b) continuing pr-training of TinyLlama-1.1B.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3.7M 24 27 22.0 0.017 0.0 0.0 0.0 0.0
32 27 24.0 0.023 0.0 0.0 0.0 0.0
32 36 28.0 0.031 93.75 76.38 87.50 65.48
9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
96 25 19.0 0.024 62.50 55.94 40.62 44.36
128 22 17.0 0.029 51.56 45.09 39.84 35.92
4.3.4 Continual Pre-training
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny" TinyLlama-1.1B [70] model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days. Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences.
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of examples. In Table 3(b), we replicate similar experiments to those in Table 2(a) for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 3(a), where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 7 in Appendix H.
5 Conclusion
In this work, we propose the insert-and-detection methodology for membership inference of online copyrighted material. Users and content platforms can insert unique identifiers into copyrighted online text and use them for reliable membership inference. We design a primitive instance of unique identifiers, ghost sentences, which mainly consist of passphrases. For demonstration purposes, we show how users can use the perplexity test and user-friendlylast-$k$ words test for the training membership detection of ghost sentences. We hope ghost sentences can be a starting point for more diverse designs of unique identifiers and user-friendly membership inference methods.
Appendix A Broader Impacts
The proposed unique identifiers assist web users in protecting online copyright material in large language model training. Ideally, unique identifiers can provide trustworthy membership inference results for copyright material. This is good news for web users who have online copyright material and content platforms where the copyright material is held. Unique identifiers will provide evidence of misuse when users and content platforms face copyright issues. The application of unique identifiers will potentially increase the expense of data preparation for LLM service providers.
Appendix B More Related Works
Instruction Tuning
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning [63]. It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks [59, 11, 45, 34, 66]. The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 [59, 45], or real data from web users [11, 34].
Diceware Passphrase
A passphrase, similar to passwords, is a sequence of words used for authentication [50]. Diceware is a method for creating passphrases by randomly selecting words from a diceware word list [53]. This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
Appendix C Verbatim Memorization Capability of Commercial LLMs
(a) Harry Potter and the Philosopher’s Stone.
(b) A Game of Thrones, Fire and Ice.
Figure 5: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5. Similar experiments and results are presented in [28].
Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 5. Some conclusions can be drawn from the phenomenon: 1) This demonstrates the significant memorization capacity of LLMs. 2) OpenAI may not have a strict process for deduplicating repeated content in the training data. Otherwise, verbatim memorization would not be possible. It is also possible that a strict deduplication process could lead to worse performance of LLMs, especially for short pieces of text, as this could break the integrity of the whole text.
Appendix D Statistics of Users on Reddit
(a) The number of documents per user.
(b) The cumulative document proportion.
Figure 6: Statistic of Reddit user data [62]. (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
Figure 6 displays the statistics of users in Webis-TLDR-17 [62], which contains Reddit subreddits posts (submissions & comments) containing "TL;DR" from 2006 to 2016. Figure 6(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 6(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data.
Table 4: Results on HellaSwag and MMLU. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12. U-Acc and D-Acc refer to Table 2(a).
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
OpenLLaMA-3Bv2 [19] 69.97 26.45
148K 256 17 13.5 2.99 71.23 26.01
128 17 13.0 1.47 71.32 26.10
64 17 13.0 0.74 71.46 26.13
32 18 12.0 0.39 71.39 26.36
16 13 11.5 0.14 71.43 25.85
16 21 16.5 0.22 70.94 25.40
8 18 13.0 0.10 71.35 26.29
8 31 25.5 0.16 71.32 25.38
4 32 32.5 0.087 71.00 25.96
2 48 47.5 0.064 70.88 25.74
1 45 45.0 0.030 70.39 25.37
1 51 51.0 0.034 70.40 25.37
148K 16 24 20.5 0.259 70.55 26.21
592K 0.065 70.76 26.64
1.8M 0.022 71.07 26.51
Appendix E Results on Common Benchmarks
In Table 4, we provide the results for instruction tuning on common benchmarks like HellaSwag [68] and MMLU [24]. Table 4 corresponds to identification results in Table 2(a). Table 4 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.